OK!
If BR2 gets 6,7 or 8 on first Hit he Stands:
With total 21 he has 75.53%(d = 17,18,19,20,21), with total 20 - 68.59%(d=17,18,19,20), with total 19 - 61.65%(d=17,18,19)
So, for these cases we have 1/13 x (75.53 + 68.59 + 61.65) = 15.83%
Now we have to consider 9 or T on first hit (Hard 12 and 13):
On Hard 12 (1/13) if we draw 6,7,8,9 (1/13 each) we should Stand and our chances to advance are 1/13 x (75.53 + 68.59 + 61.65 + 48.79) PLUS
on A,2,3,4,5 (1/13 each) we should Hit again and our chances to advance are NO BUST (Lucky Charlie) 1/13 x (4/13 + 5/13 + 6/13 + 7/13 + 8/13)
That is 19.58% + 17.75% = 37.33% , so Probability to advance for this case is 1/13 x 37.33% = 2.87%
On Hard 13 (4/13) if we draw 5,6,7,8 (1/13 each) we should Stand and our chances to advance are same as above (19.58%) PLUS
on A,2,3,4 (1/13 each) we should Hit again and our chances to advance are 1/13 x (7/13 + 6/13 + 5/13 + 4/13) = 13.02%
That is 19.58% + 13.02% = 32.6%, so Probability to advance for this case is 4/13 x 32.6% = 10.03%
The last group is A,2,3,4,5 (1/13 each) on first Hit for Soft 14 through 18.
Here on second Hit (4 cards total) we can lock Lucky Charlie (5 cards no bust, pays 3:1) and automatically advance if we will total soft 21 or less.
Thus for Soft 14 (draw A) we have 7/13, for Soft 15 (draw 2) we have 6/13, for Soft 16 (draw 3) we have 5/13, for Soft 17 (draw 4) we have 4/13, and for Soft 18 (draw 5) we have 3/13 to advance with 100% of Probability.
So for A on first Hit we have 1/13 x (7/13 x 100% + 1/13 x 9/13 + 1/13 x 8/13 + 4/13 x 7/13) = 6.19%
Here the first addendum is to have after second Hit Soft 15 through Soft 21, second addendum is to get 8 on second Hit and then hit again and Not bust, third addendum is to get 9 and Not bust on 5th card, fourth addendum is to get T and Not bust on 5th card.
Same logic will apply to Soft 15 (draw 2), Soft 16 (draw 3), Soft 17 (draw 4), Soft 18 (draw 5)
We have accordingly:
Soft 15 1/13 x (6/13 x 100% + 1/13 x 9/13 + 1/13 x 8/13 + 1/13 x 7/13 + 4/13 x 6/13) = 5.74%
Soft 16 1/13 x (5/13 x 100% + 1/13 x 9/13 + 1/13 x 8/13 + 1/13 x 7/13 + 1/13 x 6/15 + 4/13 x 5/13) = 5.28%
Soft 17 1/13 x (4/13 x 100% + 1/13 x 9/13 + 1/13 x 8/13 + 1/13 x 7/13 + 1/13 x 6/13 + 1/13 x 5/13 + 4/13 x 4/13) = 4.68%
Soft 18 1/13 x (3/13 x 100% + 1/13 x 9/13 + 1/13 x 8/13 + 1/13 x 7/13 + 1/13 x 6/13 + 1/13 x 5/13 + 1/13 x 4/13 + 4/13 x 48.79%) = 4.70% (last addendum indicates Stand on four cards total 18)
Now we can Total: 15.83% (6,7,8 on first Hit) + 2.87% (9 on first Hit) + 10.03% (T on first Hit) + 6.19% (A) + 5.74% (2) + 5.28% (3) + 4.68% (4) + 4.7% (5) = 55.32%
Previously I got 48.6%, but actually that was an estimate (some shortcuts considering not optimal play)
As we can see Hit initial Soft 13 to Soft 19 or Hard 18 is way better (55.32% vs less than 44.05) than Double Down, having Lucky Charlie option handy.